the momentum of the bed is 300 kg, m/s with the velocity of 20 m/s, what is its mass?​


1
Expert's answer
2022-05-22T18:03:23-0400
"p=mv=>m=\dfrac{p}{v}"

"m=\dfrac{300kg\cdot m/s}{20m/s}=15kg"